# A central limit theorem for a generalization of the Ewens measure to random tuples of commuting permutations

### v1

@@ 4.3. Putting the estimates together and deriving the $\mathcal{J}$ integral asymptotics
With $t_{n}$ and $x_{n}$ as in (19) and (21), defined the new sequence
@@
$\int_{\mathcal{R}_{\rm min}}e^{-q_{\ell,n}(x_{n},t_{n},\theta)}\ \frac{{\rm d}\theta}{2\pi}=\frac{\lambda_{n}}{2\pi}\int_{\mathbb{R}}f_{n}(\Theta)\ {\rm d}\Theta$
@@
${\rm Err}:=-x_{n}\sum_{\delta_{1},\ldots,\delta_{\ell}}^{\infty}\delta_{1}^{\ell-2}\delta_{2}^{\ell-3}\cdots\delta_{\ell-2}\delta_{\ell-1}^{0}\delta_{\ell}^{-1}\ e^{-\delta_{1}\cdots\delta_{\ell}t_{n}}\ R(i\lambda_{n}\delta_{1}\cdots\delta_{\ell}\Theta)\ .$
@@
$|{\rm Err}|\leq x_{n}\sum_{\delta_{1},\ldots,\delta_{\ell}}^{\infty}\delta_{1}^{\ell-2}\delta_{2}^{\ell-3}\cdots\delta_{\ell-2}\delta_{\ell-1}^{0}\delta_{\ell}^{-1}\ e^{-\delta_{1}\cdots\delta_{\ell}t_{n}}\times\frac{|i\lambda_{n}\delta_{1}\cdots\delta_{\ell}\Theta|^{3}}{6}\ ,$
@@ 5. Completion of the proofs of the main proposition and theorem
$\Psi_{\ell,n}(s)$ $=$ $-\frac{sa_{n}}{b_{n}}+\left[nt_{n}+x_{n}Z_{\ell-1}^{[\ell]}(t_{n})\right]-\left[nt_{n}+xZ_{\ell-1}^{[\ell]}(t_{n})\right]+R_{\ell,n}(s)$ (35)
@@
$a_{n}$ ::=:= $xZ_{\ell-1}^{[\ell]}(t_{n})$
$b_{n}$ ::=:= $\frac{(x\mathcal{K}_{\ell})^{\frac{1}{2\ell}}}{\sqrt{\ell(\ell-1)}}\times n^{\frac{\ell-1}{2\ell}}\ ,$
@@
$a_{n}$ ::=:= $\mathbb{E}\mathsf{K_{\ell,n}}$
$b_{n}$ ::=:= $\sqrt{{\rm Var}(\mathsf{K}_{\ell,n})}\ ,$
@@
Acknowledgements: A. A. thanks Bernhard Heim and Markus Neuhauser for instructive discussions related to the log-concavity of $\mathcal{H}_{\ell,n}(1)$ and related Bessenrodt-Ono inequalities. A. A. also thanks Ofir Gorodetsky and Mark Wildon for enlighting comments or posts on MathOverflow related to the asymptotics of $Z_{1,0}^{[2]}(t)$. S. S. gratefully acknowledges a Simons collaboration grant. S. S. also thanks Jim Pitman for pointing out useful references.

### v2

@@ 4.3. Putting the estimates together and deriving the $\mathcal{J}$ integral asymptotics
With $t_{n}$ and $x_{n}$ as in (19) and (21), define the new sequence
@@
$\int_{\mathcal{R}_{\rm maj}}e^{-q_{\ell,n}(x_{n},t_{n},\theta)}\ \frac{{\rm d}\theta}{2\pi}=\frac{\lambda_{n}}{2\pi}\int_{\mathbb{R}}f_{n}(\Theta)\ {\rm d}\Theta$
@@
${\rm Err}:=-x_{n}\sum_{\delta_{1},\ldots,\delta_{\ell}=1}^{\infty}\delta_{1}^{\ell-2}\delta_{2}^{\ell-3}\cdots\delta_{\ell-2}\delta_{\ell-1}^{0}\delta_{\ell}^{-1}\ e^{-\delta_{1}\cdots\delta_{\ell}t_{n}}\ R(i\lambda_{n}\delta_{1}\cdots\delta_{\ell}\Theta)\ .$
@@
$|{\rm Err}|\leq x_{n}\sum_{\delta_{1},\ldots,\delta_{\ell}=1}^{\infty}\delta_{1}^{\ell-2}\delta_{2}^{\ell-3}\cdots\delta_{\ell-2}\delta_{\ell-1}^{0}\delta_{\ell}^{-1}\ e^{-\delta_{1}\cdots\delta_{\ell}t_{n}}\times\frac{|i\lambda_{n}\delta_{1}\cdots\delta_{\ell}\Theta|^{3}}{6}\ ,$
@@ 5. Completion of the proofs of the main proposition and theorem
$\Psi_{\ell,n}(s)$ $=$ $-\frac{sa_{n}}{b_{n}}+\left[nt_{n}+x_{n}Z_{\ell-1}^{[\ell]}(t_{n})\right]-\left[nt_{n}+xZ_{\ell-1}^{[\ell]}(t_{n})\right]+R_{\ell,n}(s)\$(35)
@@
$a_{n}$ $:=$ $xZ_{\ell-1}^{[\ell]}(t_{n})$
$b_{n}$ $:=$ $\frac{(x\mathcal{K}_{\ell})^{\frac{1}{2\ell}}}{\sqrt{\ell(\ell-1)}}\times n^{\frac{\ell-1}{2\ell}}\ ,$
@@
$a_{n}$ $:=$ $\mathbb{E}\mathsf{K_{\ell,n}}$
$b_{n}$ $:=$ $\sqrt{{\rm Var}(\mathsf{K}_{\ell,n})}\ ,$
@@
Acknowledgements: A. A. thanks Bernhard Heim and Markus Neuhauser for instructive discussions related to the log-concavity of $\mathcal{H}_{\ell,n}(1)$ and related Bessenrodt-Ono inequalities. A. A. also thanks Ofir Gorodetsky and Mark Wildon for enlighting comments or posts on MathOverflow related to the asymptotics of $Z_{1,0}^{[2]}(t)$. We again thank Ofir Gorodetsky for explaining to us the connection of our work to the CLT by Maples, Nikeghbali and Zeindler and the mean asymptotics by Ercolani and Ueltschi. S. S. gratefully acknowledges a Simons collaboration grant. S. S. also thanks Jim Pitman for pointing out useful references.